40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 8 18:35:09 2019, gautam, Update, LSC, Locking prep 
    Reply  Wed Oct 9 18:15:26 2019, gautam, Update, LSC, PRMI 3f locking PRMI_1f.png
       Reply  Wed Oct 9 22:02:58 2019, gautam, Update, LSC, REFL55 whitening issue 
          Reply  Thu Oct 10 01:12:56 2019, gautam, Update, LSC, Locking studies PRFPMIstudies.png
             Reply  Wed Oct 16 11:42:17 2019, gautam, Update, LSC, Poor separation of PRCL/MICH in 3f signals sensMat.pdf
                Reply  Thu Oct 17 11:19:28 2019, gautam, Update, LSC, Locking activity last night MICH2PRCLnulling.pdfconsolidatedSensingMatrices.pdf.zip
                   Reply  Tue Oct 22 00:52:27 2019, gautam, Update, LSC, Locking updates PRMI3f_ALS_Oct21sensMat.pdf
                      Reply  Tue Oct 29 01:24:45 2019, gautam, Update, LSC, More locking updates POP22_feature.pngPRMI3f_ALS_Oct21sensMat.pdf
                         Reply  Tue Oct 29 15:13:19 2019, gautam, Update, LSC, More locking updates PRMI_lockloss.png
                            Reply  Wed Oct 30 01:27:00 2019, gautam, Update, LSC, More locking updates POP_ASS.png
                               Reply  Wed Oct 30 11:53:41 2019, gautam, Update, LSC, MICH loop shape tuning MICH_ELP.pdfMICH_ELP_TFs.pdf
                               Reply  Wed Oct 30 17:08:40 2019, gautam, Update, LSC, POP22 investigation POP22anomaly.pdf
                                  Reply  Mon Nov 4 15:29:47 2019, gautam, Update, LSC, POP signal path POPchain.pdf
                                     Reply  Mon Nov 4 16:06:58 2019, gautam, Update, LSC, POP optical path IMG_8027.JPGlensRealignment.jpgPOPrealigned.pngPOPdither.pngPRMfixed.png
                                        Reply  Tue Nov 5 11:52:27 2019, gautam, Update, LSC, Locking notes PRMI3f_ALS_Nov4sensMat.pdf
                                           Reply  Wed Nov 6 02:08:48 2019, gautam, Update, LSC, Locking updates 
                                              Reply  Wed Nov 6 17:05:45 2019, gautam, Update, LSC, CARM calibration CARMscan.pdfCARMcalib.pdfscan.hdf5.zip
                                              Reply  Wed Nov 6 17:45:34 2019, gautam, Update, LSC, ~ 
                                                 Reply  Mon Nov 18 21:04:38 2019, gautam, Update, LSC, Locking - some ideas 
                                                    Reply  Wed Nov 20 01:07:18 2019, gautam, Update, LSC, Locking - progress PRFPMI.png
                                                       Reply  Wed Nov 20 12:14:17 2019, gautam, Update, LSC, Locking - progress TRX_TRY_comparison.pdf
                                                          Reply  Wed Nov 20 17:20:24 2019, Yehonathan, Update, LSC, QPD Investigation 
                                                             Reply  Wed Nov 20 17:52:00 2019, gautam, Update, LSC, QPD MEDM screen update newLookQPD.pngTRX_TRY_comparison.pdfIMG_8186.JPG
                                                                Reply  Wed Nov 20 21:29:28 2019, gautam, Update, LSC, PRG ~13 PRG13.pdf
                                                                   Reply  Fri Nov 22 00:54:14 2019, gautam, Update, LSC, locking notes ALSnoiseIncrease.pdf
                                                                      Reply  Mon Nov 25 19:11:22 2019, gautam, Update, LSC, ALS noise re-look ALSnoise.pdf
                                                                         Reply  Wed Nov 27 23:24:01 2019, gautam, Update, LSC, No shaking but no inspiration either 
                                                                            Reply  Sun Dec 1 13:38:48 2019, rana, Update, LSC, No shaking but no inspiration either 
                                                                               Reply  Tue Dec 3 00:03:57 2019, gautam, Update, LSC, Green ALS also shows elevated noise with high arm buildup ALSnoiseIncrease_greenBeat.pdf
                                                                                  Reply  Wed Dec 4 19:54:27 2019, gautam, Update, LSC, A look to the past DRFPMI_2016March.pdf
                                                                                     Reply  Thu Dec 5 01:54:39 2019, gautam, Update, LSC, Partial CM board path engaged AOengaged.pdf
                                                                                        Reply  Fri Dec 6 00:02:48 2019, gautam, Update, LSC, What is the correct way to set the 3f offsets? 
                                     Reply  Fri Nov 15 12:08:04 2019, gautam, Update, LSC, POPDC whitening board 
                         Reply  Tue Oct 29 17:40:48 2019, gautam, Update, LSC, More locking updates broadSweep.pdfzoomSweep.pdf
Message ID: 15039     Entry time: Wed Nov 20 17:20:24 2019     In reply to: 15038     Reply to this: 15040
Author: Yehonathan 
Type: Update 
Category: LSC 
Subject: QPD Investigation 

{Gautam, Yehonathan}

In search of the source of discrepancy between the QPD readings in the X and Y arms, we look into the schematics of the QPD amplifier - DCC #D990272.

We find that there are 4 gain switches with the following gain characteristics (The 40m QPD whitening board has an additional gain of 4.5):

S4 S3 S2 S1 V/A
0 0 0 0 2e4
0 0 0 1 2e5
0 0 1 0 4e4
0 0 1 1 4e5
0 1 0 0 1e5
0 1 0 1 1e6
0 1 1 0 2e5
0 1 1 1 2e6
1     0 5e2
1     1 5e3

Switch 4 bypasses the amps controlled by switch 2 and 3 when it is set to 1 so they don't matter in this state.

Note that according to elog-13965 the switches are controlled through the QPD whitening board by a XT1111a Acromag whose normal state is 1.

Also, according to the QPD amplifier schematics, the resistor on the transimpedance, controlled by switch 1, is 25kOhm. However, according to the EPICS it is actually 5kOhm. We verify this by shining the QPD with uniform light from a flashlight and switching switch1 on and off while measuring the voltages of the different segments. The schematics should be updated on the DCC.

Surprisingly, QPDX switches where 0,0,0,0 while QPDY switches where 1,0,0,1. This explains the difference in their responses.

We check by shining a laser pointer with a known power on the different segments of QPDX that we get the expected number of counts on the ADC and that the response of the different segments is equal.
 

gautam edits:

  1. Lest there be confusion, the states of the switches in the (S1, S2, S3, S4) order are (0,0,0,0) for QPDX and (0,1,0,1) for QPDY.
  2. The Acromag XT1111 is a sinking BIO unit - so when the EPICS channel is zero, the output impedance is low and the DUT (i.e. MAX333) is shorted to ground. So, the state of the MAX333 shown on the schematics corresponds to EPICS logic level 1, and the switched state corresponds to logic level 0.
  3. For the laser pointer test, we used a red laser pointer. Using a power meter, we measured ~100uW of 632nm power. However, we think this particular laser pointer had failing batteries or something because the spot looked sometimes brighter/dimmer to the eye. Anyways, we saw ~10,000 ADC counts when illuminating a single segment (with the QPD gain switches at the 0,0,0,0 setting, before we changed anything). We expect 100uW * 0.4 A/W * 500 V/A * 10 * 40 * 4.5 * 3267.8 cts/V = ~12000 cts. So everything seems to check out. We changed the gain to the 5kohm setting and bypassed the subsequent gain stages, and saw the expected response too. The segments were only balanced to ~10%, but presumably this can be adjusted by tweaking digital gains.
ELOG V3.1.3-